OpenStudy (anonymous):

give the function f(x)=-8x+1 find f-1 (-15)

OpenStudy (anonymous):

do u mean f^-1(-15)

OpenStudy (anonymous):

Find the inverse function first by swapping x and y. x=-8y+1 f^-1(x)=-(x-1)/8 -(-15-1)/8=16/8=2

OpenStudy (anonymous):

Thank you I had difficulty finding the steps to the solution.